OpenStudy (anonymous):

In a change jar, Tom has nickels and dimes. The number of nickels is two times the number of dimes and the total amount of money in the jar is $1.20. How many dimes are in the jar?

OpenStudy (anonymous):

Nickel = 5 cents Dime = 10 cents Let number of nickels be N. Number of dimes = 1/2 (Number of Nickels) = 1/2 N Total Money = Value of Nickels + Value of Dimes 1.20 = (N)(0.05) + (1/2N)(0.1) 1.20 = 0.05N + 0.05N N = ?

OpenStudy (anonymous):

Thanks for your help